The perimeter of a rectangle is 126cm the rectangle is twice as long as it is wide. What are the dimensions
Answer by macston(5194)

W=width L=length=2W; P=perimeter
P=2(L+W)
126cm=2(L+W) Divide each side by 2.
63cm=L+W Substitute for L.
63cm=2W+W
63cm=3W Divide each side by 3.
21=W ANSWER 1: The width is 21 cm.
L=2W=@(21cm)=42cm ANSWER 2: The length is 42 cm.
CHECK:
126cm=2(L+W)
126cm=2(42cm+21cm)
126cm=2(63cm)
126cm=126cm